OpenStudy (anonymous):

The diameter of a circle drawn on a coordinate grid has one of its end points at (-3, 2) and the other end point at (6, -4). Leena performed the steps shown to find the length of the diameter. Steps are in the bottom

OpenStudy (anonymous):

In which step did Leena first make an error? Step 1, because she substituted the incorrect values for x2 and y1. Step 3, because she added the x coordinates instead of subtracting them. Step 2, because she used the incorrect formula to find the length. Step 5, because she found the square root of 45 instead of the square root of 9

OpenStudy (anonymous):

I did the formula and got 45 too so it can't be step 5. It's the proper formula so not step two, she had to add because two negatives is a positive, and she didn't substitute them wrong, so what else?

zepdrix (zepdrix):

hmm link doesn't work <:o

OpenStudy (anonymous):

Sec i'll write it out

OpenStudy (anonymous):

Step 1. x1 = -3 y1 = 2 x2 = 6 y2 = -4 Step 2. Distance \[\sqrt{(x1-x2)^2 - (y1-y2)^2 }\]

OpenStudy (anonymous):

Step 3. Distance \[\sqrt{(6+3)^2 - (-4-2)^2}\]

OpenStudy (anonymous):

Step 4. \[\sqrt{(9)^2 - (-6)^2 }\]

OpenStudy (anonymous):

Step 5. Distance is approximately 6.71 That's all of them.

zepdrix (zepdrix):

\[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] the fact that the x2 and x1 are reversed shouldn't matter, since they're being squared anyway. But i think the problem is the negative, im pretty sure the distance formula is the sum of the squares, not differences. Step 2 seems like the bad one :O

OpenStudy (anonymous):

Can't believe I missed that.

zepdrix (zepdrix):

hehe c: happens to all of us from time to time XD

OpenStudy (anonymous):

All my distance questions have been turning up wrong, I'm here frustrated because none of my answers are included in here. Thanks alot man. c: